Negative thirty-one subtracted from -19
Answer by lenny460(1073)

Negative thirty-one subtracted from -19
I am getting two answers from your question above:
First Answer:
-19 - 31 = -50
Answer: -50

Second Answer:
-19 - (-31)
-1 * -31 = +31
Therefore:
-19 + 31 = 12
Answer:
12


Lennox Obuong
Algebra Tutor